OpenStudy (anonymous):

A pair of dice is rolled together till a sum of either 5 or 7 is obtained. The probability that 5 comes before 7 is...

OpenStudy (anonymous):

@experimentX another question...

OpenStudy (anonymous):

Shall I show what i did?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

the probability that 5 comes before 7 is 4/10=2/5

OpenStudy (anonymous):

Solution please..@Shriya

OpenStudy (anonymous):

sample set= [(1,4), (2,3), (3,2), (4,1), (1,6) (2,5), (3,4), (4,3), (5,2), (6,1)] so, n(S), i.e., the total number of elements in S=10 out of this, the first four terms show the sum=5. therefore n(5)=4. therefore, P(5)=n(5)/n(s) therefore, P(5)=4/10=2/5

OpenStudy (anonymous):

No..this can't be it..the sample space does not contain the right number..it should be 36 for two dice, whatever may be the condition

OpenStudy (anonymous):

I got the same answer but in a different way...but I don't think your method is right

OpenStudy (anonymous):

ok u calculated it considering the entire sample space. i considered just the points those are required. whatever be the criteria, the ratio remains the same. n hence our answers are mathing!

OpenStudy (anonymous):

Yea..I got you now..that's right..Thanx for a new method !!

OpenStudy (anonymous):

wc :)

OpenStudy (kropot72):

There are 4 ways of getting a sum of 5 and 6 ways of getting a sum of 7. The sample space has 36 outcomes. The probability of getting a sum of 5 or a sum of 7 is \[P(5or7)=\frac{4}{36}+\frac{6}{36}=\frac{10}{36}\] \[P(5first)=\frac{P(5)}{P(5or7)}=\frac{4}{36}\times \frac{36}{10}=0.4\]

OpenStudy (anonymous):

Thank you @kropot72 for your answer..

OpenStudy (kropot72):

You're welcome :)